# Mingyu Liu1    Jiuhe Shu1∗    Hui Chen1    Zeju Li1    Canyu Zhao1 Jiange Yang2    Shenyuan Gao3    Hao Chen1    Chunhua Shen1

### v1

@@ 4.4 Action Linear Probing Experiment
We then train a lightweight Multi-Layer Perceptron (MLP) to predict the action sequence $\mathbf{A}_{n}$ from the latent action representation $\Delta z$. The performance is measured by the Mean Squared Error (MSE) between the predicted action sequence $\hat{\mathbf{A}}_{n}$ and the ground-truth sequence $\mathbf{A}_{n}$. This process can be formally described as follows:
@@
Figure 5: Scaling Performance. The Performace of our model can be steadly scaling with more data, including human ego-centric data.
To validate our approach, we benchmark our latent action representation against several baselines. The first set of baselines tests the importance of our structured latent space against alternatives operating on raw data: Pooled Delta Image (using pixel-wise differences) and Delta DINOv2 Features (using feature-space differences). The second baseline, LAPA (Ye et al., 2024b), contrasts our deterministic state-difference formulation with a state-of-the-art autoregressive model that generates latent action tokens.
For a fair comparison, all representations are used to train an identical lightweight MLP. We evaluate non-generative methods across 1, 2, 4, and 8-step prediction horizons, while LAPA is assessed at its native 1-step horizon. Once trained, the MLP can directly map a current observation and a goal image to an executable action trajectory on the test set. The results in Figure 5 confirm the superiority of our method. Additional details are provided in Appendix B.
@@ 4.5 Real World Experiments
Experiment Setup. Our experimental benchmark comprises six tasks—three short-horizon and three long-horizon—to comprehensively evaluate the effectiveness of our representation in facilitating world modeling and decision-making across a spectrum of task complexities, as shown in Figure 6. For each task, we collect 50 demonstrations and evaluate performance over 20 trials, results are shown in Table  5, our real-world robot setting and tasks definition are detailed in Appendix C.
Table 5: Performance comparison on short and long horizon tasks.
@@
OpenVLA 0.35 0.3 0.25 0.30 0.2 0.3 0.15 0.2 0.25
OpenVLA + StaMo states(S) 0.5 0.45 0.3 0.42 0.25 0.3 0.2 0.25 0.34
OpenVLA + StaMo states(S+D) 0.6 0.6 0.55 0.58 0.45 0.6 0.45 0.5 0.53
OpenVLA + StaMo states(S+D+O) 0.65 0.65 0.5 0.6 5 0.5 0.45 0.52 5.56
OpenVLA + StaMo states(S+D+O+Ego) 0.7 0.75 0.6 0.72 0.65 0.65 0.55 0.62 0.67
4.6 Scalable Simulation and Real World Performance
A natural question following our initial validation is the scalability of our approach. To explore this, we scaled our training with a diverse corpus of data, including more simulation, the Open X-Embodiment dataset(which contains more different robot embodiments), and egocentric human demonstrations. The results in Table 5 reveal steady performance improvements across both simulated and real-world robotics benchmarks, which clearly depicted in Figure 5, confirming that our model is highly scalable and benefits from diverse data sources.
5 Conclusion
In this work, we introduced StaMo, a framework that extracts reusable latent action vectors from static images via a transformer-based tokenizer and a diffusion decoder. We demonstrate that linear interpolation in this latent space produces smooth, physically plausible motions, and that these vectors generalize zero-shot to unseen robots—suggesting that large-scale visual models implicitly learn a linearized dynamics manifold. Our approach offers a pathway toward scalable unsupervised skill discovery from diverse visual data, bridging the gap between static perception and dynamic action.
Ethic Statement
This research adheres to widely recognized ethical guidelines in robotics and artificial intelligence. All experiments were conducted using publicly available datasets and simulated environments, without involving personal data, or sensitive information. The methods proposed in this work are intended solely for academic research and technological advancement, and do not pose immediate risks of misuse. We encourage the responsible deployment of robotic technologies, ensuring that their applications align with societal values and ethical standards.
Repoducibility Statement
To ensure reproducibility, we provide detailed descriptions of the model architectures, training procedures, and evaluation protocols in the Section 3. All experiments were conducted using publicly available datasets (e.g., LIBERO and DROID) and standard simulation environments. Hyperparameters and implementation details are fully documented in Appendix A. The code, configuration files, and instructions necessary to reproduce our results will be released upon publication.
Appendix A Implementation Details
Our StaMo encoder experiments were conducted on Ubuntu 22.04 LTS with PyTorch 2.1, using publicly available pretrained weights such as DinoV2 and Stable Diffusion 3 for model initialization. Unless otherwise noted, all main experiments were carried out on standard public datasets under a consistent hardware and software setup. Training was performed with the AdamW optimizer, an initial learning rate of 3e-5 with a cosine decay scheduler, a batch size of 512 per GPU, and a weight decay of 1e-3. For evaluation, PSNR and SSIM were adopted as the primary metrics to assess the quality of image reconstruction. All experiments were run on 8 NVIDIA H100 GPUs, with a typical training duration of around ten days, and we fixed the random seed to 33 to ensure reproducibility.
Unless otherwise specified, all main experiments are conducted using data from publicly available datasets, LIBERO and DROID. We strictly adhere to their original release protocols and only perform essential preprocessing steps, without involving any private or sensitive information.
Figure 7: Visual reconstruction results from the same episode. The first and last frames are reconstructed from ground-truth images using the StaMo encoder, while the intermediate frames are generated by linearly interpolating between the latent state tokens of the two endpoints. The transitions show that both the robotic arm and the objects move in a continuous and smooth manner.
Figure 8: Transfer linear interpolation experiment with the StaMo encoder. The left and right panels illustrate different task scenarios, where reconstructions are obtained by tokens(3) + tokens(2) – tokens(1), demonstrating the linear interpolation property of latent representations during transfer.
@@ Appendix A Implementation Details
As shown in Table 6, all results are obtained from models trained on the libero datasets, where the latent state is represented using only 2 tokens. Across different hidden dimensions (256/512/1024), the model achieves comparable PSNR and SSIM scores on all datasets, with only marginal variations. This indicates that the choice of hidden dimension has limited impact on reconstruction performance, and the model maintains stable effectiveness across a wide range of dimensional settings.
Appendix B Action Linear Probing Details
Our linear probing dataset is constructed by sampling from the LIBERO dataset. We performed a randomized sampling process across all available trajectories to create a dataset of tuples $(\mathbf{I}_{n},\mathbf{I}_{n+k},\mathbf{A}_{n})$, as described in the main paper. A total of approximately 20000 samples are collected, which are then partitioned into training and validation sets with a 95%/5% split, respectively. This process is repeated independently for each of our four evaluated action horizons, where $k\in\{1,2,4,8\}$. To ensure a fair comparison, the same random seed are used for data sampling and splitting. Notably, to accommodate the LAPA baseline which requires task description input, each sampled tuple also includes the corresponding language instruction. For the $k=1$ comparison, the exact same initial frames ($\mathbf{I}_{n}$) are maintained through using same random seed across all methods.
Following the standard linear probing protocol, we use a simple Multi-Layer Perceptron (MLP) as the probe to decode the action representations. The MLP consists of a single hidden layer with 128 units and a ReLU activation function. The architecture is intentionally kept minimal to ensure that the measured performance primarily reflects the quality of the input representation rather than the capacity of the probe itself. During training, the upstream encoders (i.e., StaMo and DINOv2) are kept frozen, and gradients are only computed and backpropagated for the MLP’s parameters.
To visually substantiate the rich motion information encoded within our latent actions, we conduct a qualitative replay experiment in the LIBERO simulation environment. For each of the ten tasks in the LIBERO-10 benchmark, we utilize all the ground-truth initial frames ($\mathbf{I}_{n}$) and final frames ($\mathbf{I}_{n+k}$) from the demonstrations. These two frames are encoded into a single latent actions ($\Delta z$) using our frozen StaMo encoder. Here we use $k=8$ in our experiment.
The latent actions are then fed into the corresponding pretrained MLP probe to predict the entire sequence of $k$ actions. To lower the accumulation of small prediction errors, we only execute the first action in every sequences. The results, visualized in Figure 9, show that even the actions decoded by this simple MLP are sufficient to successfully execute the corresponding tasks to some degree. We also observe that the primary failure mode occurred in tasks requiring precise grasping, where the predicted actions sometimes cause the gripper to descend to an insufficient depth.
Figure 9: MLP-predicted actions replay. Actions are predicted by an MLP probe from our StaMo latent action representation and replayed in LIBERO-10 tasks. The results show that our latent actions are remarkably effective, encoding functionally coherent and executable motions.
Appendix C Real World Setting
Our robotic setup includes a single Franka Research3 arm equipped with a UMI (Chi et al., 2024) grippper. A third-person-view RealSense D435 camera is mounted in a fixed position to capture environmental observations at a resolution of 640×480 pixels. Following the implementation in publicly available code  [^1] https://github.com/UT-Austin-RPL/deoxys_control, we use a 3D mouse for teleoperated data collection. Our system operates at 20 Hz, which is a moderate reduction from the native 100 Hz control frequency to balance training efficiency and motion continuity. Actions are defined as absolute end effector poses in SE(3), using 3D position and quaternion orientation. Our task definition are list below:
1) Pick up the [TOY NAME]: The robot is required to pick up the specific toy, including the toy snake, the toy eggplant, and the toy chicken, resulting in a total of three tasks (with no distractions).
2) Put the toy into the basket: The robot need to pick up the toy(no other distractions) and put it into the basket.
3) Open the drawer: The robot is required to approach the drawer and pull it open.
4) Put all cups into the basket: We place two cups with different colors on the table and set a few other things(toys) as distractions. The robot need to sequentially put two cups into the basket.
5) Put the toy into the drawer: We place a toy on the table and set a few other things as distractions. The robot need to sequentially open the drawer, pick up the specific toy and put it into the drawer, and close the drawer.
6) Stack all cups: Three cups of varying sizes are placed on a tabletop. The robot’s task is to perceive their relative sizes and then stack them in descending order (from largest to smallest).drawer, and close the drawer.

### v2

@@ 4.6 Action Linear Probing Results
We then train a lightweight Multi-Layer Perceptron (MLP) to predict the action sequence $\mathbf{A}_{n}$ from the latent action representation $\Delta z$. The performance is measured by the Mean Squared Error (MSE) between the predicted action sequence $\hat{\mathbf{A}}_{n}$ and the ground-truth sequence $\mathbf{A}_{n}$. This process can be formally described as follows.
@@
To validate our approach, we benchmark our latent action representation against several baselines. The first set of baselines tests the importance of our structured latent space against alternatives operating on raw data: Pooled Delta Image (using pixel-wise differences) and Delta DINOv2 Features (using feature-space differences). The second baseline, LAPA [54], contrasts our deterministic state-difference formulation with a state-of-the-art autoregressive model that generates latent action tokens.
For a fair comparison, all representations are used to train an identical lightweight MLP. We evaluate non-generative methods across 1, 2, 4, and 8-step prediction horizons, while LAPA is assessed at its native 1-step horizon. Once trained, the MLP can directly map a current observation and a goal image to an executable action trajectory on the test set. The results in Figure 4 confirm the superiority of our method.
4.7 Goal-Conditioned Task Planning
Once the linear head is trained, it becomes capable of translating abstract representations into concrete, executable actions. We evaluated the goal-conditioned task planning capabilities of StaMo and DINO-WM [57] using the ManiSkill benchmark. ManiSkill was specifically selected because it was not included in the pre-training datasets for either StaMo or DINO-WM. For this comparison, we trained only the linear head while keeping the encoders of both StaMo and DINO-WM frozen. Table 6 presents the performance comparison, and further implementation details are available in the appendix. We evaluated each task over 100 trials, with results presented as percentage success rates.
Table 6: Goal-Conditioned Task Planning.
Method Pick Cube Push Cube Stack Cube Pull Cube Tool
DINO-WM [57] 12% 8% 4% 18%
StaMo 22% 32% 16% 38%
@@ 4.8 Real World Experiments
Table 7: Performance comparison on short and long horizon tasks.Best results are in bold and second-best results are underlined.
@@
Spatial-VLA [39] 0.45 0.50 0.20 0.38 0.35 0.20 0.20 0.25 0.32
WorldVLA [12] 0.55 0.20 0.15 0.30 0.25 0.40 0.15 0.26 0.28
UniVLA [46] 0.55 0.30 0.35 0.40 0.35 0.30 0.10 0.25 0.33
OpenVLA [25] 0.35 0.30 0.25 0.30 0.20 0.30 0.15 0.20 0.25
OpenVLA + StaMo state 0.60 0.55 0.50 0.60 0.55 0.55 0.45 0.52 0.56
OpenVLA-OFT + StaMo state 0.70 0.65 0.60 0.65 0.70 0.65 0.55 0.63 0.64
Experiment Setup Our real-world experimental benchmark comprises six manipulation tasks (three short-horizon and three long-horizon) to comprehensively evaluate the effectiveness of our StaMo representation in facilitating world modeling and decision-making across a spectrum of task complexities, as shown in Figure 5. For each task, we collected 50 human demonstrations and evaluated performance over 20 trials. The results, presented as percentage success rates, are shown in Table 7.
4.9 Ablation Study
4.9.1 Ablation of Token Dimension
We ablate the hidden dimension of StaMo’s 2-token latent state (trained on LIBERO datasets, evaluated via PSNR/SSIM). As shown in Table 8, across 256/512/1024 dimensions. This indicates hidden dimension has limited impact on performance, as the pre-trained DiT decoder’s generative prior offsets potential information loss in lower dimensions.
@@ 4.9.1 Ablation of Token Dimension
4.9.2 Ablation of Additional Training Data
A natural question following our initial validation is the scalability of our approach. To explore this, we investigated the effect of the StaMo training data volume and diversity on real-world task performance. We compared several data configurations: (1) simulation data only, (2) simulation + DROID [23] data (our main setting), (3) simulation + DROID + OXE [35] data, and (4) simulation + DROID + OXE + egocentric human (Ego) data [44]. The results, clearly depicted in Figure 6, confirm that our model is highly scalable. This verifies that incorporating larger volumes of data and greater task diversity, such as varied robot embodiments from OXE and human-centric data, steadily improves StaMo’s performance capabilities.
Figure 6: Scaling Performance. Performance of our model can be scale with more data, including human ego-centric data.“S” means Simulation data only,“D” represents DROID [23] data, “O” represents OXE [35] data, “Ego” means Ego-centric data.
4.9.3 Ablation of Pretrained Diffusion Decoder
We compared the performance of using a pre-trained diffusion decoder against training one entirely from scratch. The decoder trained from scratch required a significantly longer time to converge and ultimately was unable to match the same level of performance as its pre-trained counterpart. As shown in Table 9. This finding demonstrates the critical importance of pre-training on large-scale natural image datasets. Such pre-training provides a crucial prior, which is highly beneficial for learning a more effective representation of the robotics data distribution.
Table 9: Performance comparison between using pretrained diffusion decoder and training from scratch.
Method Spatial Object Goal Long Average
Train from Scratch 90.3% 92.7% 86.5% 73.4% 85.7%
w. Pretrain Diffusion Decoder 92.3% 94.8% 88.1% 75.2% 87.6%
5 Conclusion
In this work, we introduced StaMo, a framework that extracts reusable latent action vectors from static images via a transformer-based tokenizer and a diffusion decoder. We demonstrate that linear interpolation in this latent space produces smooth, physically plausible motions, and that these vectors generalize zero-shot to unseen senarios, suggesting that large-scale visual models implicitly learn a linearized dynamics manifold. Our approach offers a pathway toward scalable unsupervised skill discovery from diverse visual data, bridging the gap between static perception and dynamic action.
Acknowledgments
This work was supported in part by the Zhejiang Provincial Pioneering Science and Technology Program (2025), and by the National Natural Science Foundation of China (No. 62576315).